Skip to main content

IT Status Alerts

All Closed Alerts

Go to Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 218 219 220 221 222 223 224 225 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244 245 246 247 248 249 250 251 252 253 254 255 256 257 258 259 260 261 262 263 264 265 266 267 268 269 270 271 272 273 274 275 276 277 278 279 280

5596 results found, displaying results 201 - 221.

Performance Issue: Alumni Email to AOL and Yahoo

Start:
2019-04-01 11:03:00
Status:
Closed
Brief Description:
Delays have been observed delivering alumni email to AOL and Yahoo. Cornell staff are temporarily rerouting mail to address the issue while working on a more permanent fix.
Current Status:
The issue has been resolved and CIT is continuing to monitor the situation.
more details

Unplanned Outage: Power failed to return when work was finished

Start:
2019-04-02 04:00:00
Status:
Closed
Brief Description:
Work was completed on schedule, but when power was restored, some devices did not come back on-line. Network Engineers are on-site working on Managed Firewalls to fully restore service.
Current Status:
North Campus wireless networks have been restored. VPN Service to North Campus should be restored soon. 04:40 UPDATE Eric called back to say that all services have been restored at this time. We will continue to monitor the situation until we are sure that all devices are working properly.
more details

Scheduled Service Change: Blackboard Data Center Maintenance Impacting service use - March 31, 2019

Start:
2019-03-31 02:00:00
Status:
Closed
Brief Description:
On Sunday, March 31, 2019, from 2:00am-7:00am EST, Blackboard will be performing scheduled datacenter maintenance to increase the capacity of our firewall from 1G to 10G.
more details

Unplanned Outage: Wired and Wi-Fi data network outage at HyroElectric Plant

Start:
2019-03-27 18:32:00
Status:
Closed
Brief Description:
Wired and Wi-Fi data network outage at HydroElectric Plant. Due to power outage.
Current Status:
Power to HydroElectric plant building has been restored. Network service remains unavailable at this time. CIT network technicians will investigate.
more details

Unplanned Outage: Wired and Wi-Fi data network outage at Chilled Water Plant

Start:
2019-03-28 11:03:00
Status:
Closed
Brief Description:
The cwp2-1-2540 networking switch at the Chilled Water Plant is currently down.
Current Status:
This switch is now back up and connectivity has been restored to the Chilled Water Plant.
more details

Scheduled Service Change: PeopleSoft Migration of Batch and Online Objects - March 28, 2019

Start:
2019-03-28 05:00:00
Status:
Closed
Brief Description:
PeopleSoft migration of batch and online objects to production will take place on Thursday, March 28, 2019.
more details

Scheduled Service Change: EZ-Backup tape library firmware updates - March 27, 2019

Start:
2019-03-27 10:00:00
Status:
Closed
Brief Description:
EZ-Backup team will be updating tape library firmware on Wednesday, March 27, 2019, between 10:00am-6:00pm.
more details

Scheduled Service Change: Academic Web Hosting Servers OS Conversion - March 26, 2019

Start:
2019-03-26 04:00:00
Status:
Closed
Brief Description:
Academic web hosting production servers will be converted from RedHat 6 to CentOS 6 on Tuesday, March 26, 2019, between 4:00am - 8:00am.
more details

Unplanned Outage: Wired and Wi-Fi data network outage at Seneca Place, and Sage House

Start:
2019-03-25 10:40:00
Status:
Closed
Brief Description:
The Seneca Place, and Sage House switches is currently down due to NYSEG power outage. As a result wired and Wi-Fi data network service is unavailable in those buildings.
Current Status:
Power has been restored and networking switches are now back up.
more details

Performance Issue: Video On Demand - Adding New Users as Collaborators or Channel Members

Start:
2019-03-12 14:33:00
Status:
Closed
Brief Description:
Users are unable to add other users created after 1/8/19 as collaborators or channel members in Video On Demand (VOD). CIT is working with the vendor to fix the issue.
Current Status:
The vendor has implemented a solution and new user accounts are now successfully populating in Video On Demand (VOD).
more details

Scheduled Service Change: KFS Patch Release - March 24, 2019

Start:
2019-03-24 06:00:00
Status:
Closed
Brief Description:
KFS patch release will be applied on Sunday, March 24, 2019, between 6:00am - 12:00pm.
more details

Scheduled Service Change: PeopleSoft Server OS Patching, Windows and Linux - March 24, 2019

Start:
2019-03-24 01:00:00
Status:
Closed
Brief Description:
PeopleSoft Windows and Linux Patches will be applied to servers utilized by the PeopleSoft Campus Solutions System on Sunday, March 24, 2019, between 1:00am - 9:00am.
more details

Performance Issue: SUNYNet Access Issues

Start:
2019-03-19 13:41:00
Status:
Closed
Brief Description:
CIT's network engineering department is investigating an issue accessing SUNYNet websites from Cornell networks.
Current Status:
SUNYNet services are functioning properly when accessed from Cornell. SUNY is still working on a more permanent fix, but a workaround is in place to restore services and access.
more details

Scheduled Service Change: Blackboard managed hosting maintenance, impacting service use - 3/22/2019

Start:
2019-03-22 02:00:00
Status:
Closed
Brief Description:
On Friday, March 22, 2019, between 2:00am - 6:00am ET, the Blackboard managed hosting engineers will perform service-impacting maintenance, replacing firewall devices to increase capacity.
more details

Scheduled Service Change: PeopleSoft Migration of Batch and Online Objects, 3/21/2019

Start:
2019-03-21 05:00:00
Status:
Closed
Brief Description:
PeopleSoft migration of batch and online objects to Production, 3/21/2019, 5:00am-7:00am
more details

Scheduled Service Change: EZ-Backup team to update library firmware and control path drives - 3/20/19

Start:
2019-03-20 11:00:00
Status:
Closed
Brief Description:
EZ-Backup team to update tape library firmware and control path drives - March 20, 2019, 11:00am-6:00pm
more details

Unplanned Outage: Wired and Wi-Fi data network outage in Post Lab

Start:
2019-03-20 10:03:00
Status:
Closed
Brief Description:
The postlab2-3500 switch is currently down. As a result wired and Wi-Fi data network service is unavailable at Post Lab building.
Current Status:
This switch is now back up and connectivity has ben restored to the Post Lab.
more details

Security Alert: Google Chrome critical vulnerability

Start:
2019-03-06 16:33:00
Status:
Closed
Brief Description:
Google has announced a critical vulnerability in the Chrome Browser. Cornell IT is patching for this vulnerability this week, or you can proactively update your browser.
more details

Scheduled Service Change: Monthly Updates to Net-Print PaperCut MS OS, 3/19/2019

Start:
2019-03-19 05:00:00
Status:
Closed
Brief Description:
Net-Print PaperCut: planned outage for MS OS Monthly updates, 3/19/2019, 5:00-7:00am
more details

Scheduled Service Change: Upgrade the main elist server (Y1) to Aurea LM v12.3.5 - 3/17/2019

Start:
2019-03-17 07:00:00
Status:
Closed
Brief Description:
On March 17, 2019, between 7:00am - 7:30am, we will upgrade the main e-list server (Y1) to the latest version of Aurea Listmanager 12.3.5 (from v12.3.4).
more details
Go to Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 218 219 220 221 222 223 224 225 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244 245 246 247 248 249 250 251 252 253 254 255 256 257 258 259 260 261 262 263 264 265 266 267 268 269 270 271 272 273 274 275 276 277 278 279 280